# True Zero-Shot Inference of Dynamical Systems Preserving Long-Term Statistics

### v2

@@ 1 Introduction
Figure 1: DynaMix achieves zero-shot DSR of attractor geometry and long-term temporal properties (power spectrum) from a short context signal while Chronos [2] fails.
Most real-world processes, from atmospheric phenomena and stock markets to brain activity or ecological networks, can be described as dynamical systems (DS) [56, 15, 86, 59]. Reconstructing these from observational data, called dynamical systems reconstruction (DSR), has been a long-standing challenge in scientific modeling [98, 23, 31]. DSR goes beyond conventional time series (TS) modeling, as we wish to have a generative model of the underlying process which exhibits long-term behavior with the same temporal and geometrical signatures (Fig. 1), i.e. the same invariant or ‘climate’ statistics [66, 70], as important in scientific applications. Achieving this usually requires special control-theoretic training techniques [58, 39] or loss objectives [69, 70, 42, 76] which accentuate the system’s long-term dynamics. Numerous deep learning approaches for DSR, based on recurrent neural networks (RNNs; [85, 22, 88, 16, 6, 73, 39, 7, 70]), neural ODEs [17, 45, 1, 47], Koopman operators [14, 55, 63, 3, 60, 28, 91], or library-based methods [12, 53, 43, 19, 57], have been advanced over the years. However, all of these require purpose-training on the specific system observed and struggle to generalize beyond their training distribution [33].
Inspired by the strong in-context and zero-shot generalization abilities of LLMs [10, 27, 21, 18], there has been a push recently to develop models with likewise properties for the time series domain. Time series foundation models like Chronos [2] or Mamba4Cast [5] are pre-trained on a huge database of real-world and artificial time series, and then tasked to forecast novel time series from which snippets are presented in-context, without any parameter fine-tuning. They are not built for DSR, however, and – as we show here – typically fail to properly capture a system’s long-term behavior and the structure of its attractors (Fig. 1; [99]).
@@
  • Accurate zero-shot DSR: DynaMix achieves strong generalization across diverse DS, eliminating the need for fine-tuning while maintaining accuracy in reproducing attractor geometry and long-term statistics.
@@
  • Computational and parameter efficiency: The model achieves high performance with a very lightweight architecture ($\approx 10$k parameters) and small training corpus ($34$ systems), enabling orders of magnitude faster inference than other foundation models.
@@ 2 Related work
In DSR we seek to learn generative models from time series data that represent the underlying system dynamics with all its topological, geometrical, and temporal properties [23, 33, 70, 30, 13, 11], in that sense providing an approximation to the underlying system’s governing equations. By definition, such a model should not only render viable short-term forecasts, but also reproduce the long-term evolution of the DS it has been trained on, both in state space and in the time domain. Methods approaching this goal have been founded on predefined function libraries, such as Sparse Identification of Nonlinear Dynamics (SINDy; [12, 53, 19, 43, 57]), on reservoir computers [67, 69, 70], neural ODEs [17, 45, 1, 47], Koopman operators [14, 55, 63, 3, 60, 28, 91], or different types of RNNs [85, 22, 88, 16, 6, 73, 39, 7]. More important than the architecture itself seems to be their proper training in order to ensure long-term (invariant) statistics of the underlying system are met: Methods like sparse [58, 6] or generalized [39] teacher forcing allow RNN-generated trajectories to ‘explore the future’ whilst training, yet keep loss gradients in check. Alternatively, long-term statistics based on the observed system’s Lyapunov spectrum, fractal geometry, or invariant measures may be added to regularize the loss [69, 70, 42, 76], but require to compute these properties from the data first. Despite these advances, out-of-domain generalization remains a key challenge in DSR [33]. Meta- and hierarchical learning models, trained across many DS simultaneously, have recently been designed as steps toward foundation models for DSR [9, 62, 46, 97], but still require parameter fine-tuning and lack in-context inference capabilities.
@@
Large language models (LLMs) exhibit an impressive ability to infer patterns from prompts (context) and generalize to novel situations without retraining [10, 27, 21, 18], although the underlying reasons for this are still a matter of debate [51, 90, 96]. This inspired the development of general-purpose time series (TS) foundation models which could accurately forecast time series from a short segment provided ‘in-context’ [2, 26], without the need of task-specific fine-tuning [20]. One idea is to simply use pretrained LLMs directly as TS forecasters [34, 71, 79, 93, 83], but the inherent differences between textual and many temporal data, such as continuity, present significant challenges [82]. To overcome these, transformer-based architectures, similar in design to LLMs, such as Chronos [2] or TimesFM [20], were specifically pretrained on a large corpus of time series data. Promising zero-shot forecasting capabilities have also been achieved with alternative architectural designs, such as Tiny Time Mixers [26] or state-space models like Mamba [5].
The success of TS foundation models raised hope these could also be utilized for zero-shot DSR, but the – to our knowledge – so far only previous study on this, based on Chronos, had mixed outcomes and fell short of a full DSR evaluation [99]: Successful DSR, if present at all, heavily depended on the initial conditions, and mechanisms such as context parroting visually gave the wrong illusion that features of the dynamics had been captured. In fact, here we show that existing TS foundation models are generally not capable of producing valid DSRs (Fig. 1), and highlight some of their failure modes. This is in contrast to the zero-shot DSR foundation architecture we develop here, based on models [7] and training algorithms [58] successful in DSR.
@@ 3.1 Model architecture
To enable zero-shot reconstruction of novel DS, we develop a specific mixture-of-experts (MoE) architecture that can be pretrained across many diverse DS (Fig. 10), with different experts possibly specializing in different dynamical regimes. As a SOTA DSR base model for the experts, we leverage a recent parameter-friendly RNN, which allows for highly efficient DSR training and is designed to yield topologically parsimonious and interpretable representations of DS it is being trained on, the Almost-Linear RNN (AL-RNN; [7]):
@@
$\bm{w}^{att}_{t}=\sigma\left(\frac{\left|\bm{C}-\left(\bm{Dz}_{t}+\bm{\epsilon}\right)\bm{1}_{T_{C}}^{\top}\right|^{\top}\bm{1}_{N}}{\tau_{\text{att}}}\right)\in\mathbb{R}^{T}\;,$ (3)
@@ 3.2 Model training
For training, a collection of just $34$ different chaotic and cyclic DS is used [29], see Appx. A.2 for full specification and Fig. 10 for illustration. From each, multivariate time series $\bm{X}\in\mathbb{R}^{N\times T}$ are simulated ($\approx 6\times 10^{5}$ in total) and then standardized dimension-wise (likewise for the test set), of which the first $T_{C}<T$ column entries are defined as the context signal $\bm{C}=\bm{X}_{1:T_{C}}$. The experts are initialized at $t_{0}=T_{C}-\Delta t+1$ and forward-iterated until time $T$, such that they overlap for a period $\Delta t$ with the context. This is to ensure that the model learns to optimally utilize the context for generalization. A MSE loss is then computed across ground truth, $\bm{X}_{T_{C}-\Delta t+1:T}$, and respective model-generated time series $\hat{\bm{X}}$. The model is trained by sparse teacher forcing (STF; [6, 58]), a control-theoretic technique specifically designed for DSR. STF replaces part of the forward-iterated latent states $\bm{z}_{t}$ by data-inferred states $\hat{\bm{z}}_{t}$ at optimally chosen intervals $\tau$, to avoid exploding gradients even for chaotic systems while enabling the DSR model “to explore the future” (see [58] for detailed theoretical motivation). As shown in Fig. 26, STF with optimal $\tau$ is essential for achieving good zero-shot DSR results. Note that STF is only used for training and turned off at test time.
@@ 3.3 DSR evaluation & time series forecasting
We used two established measures to assess DSR quality [48, 6, 39, 94, 99, 70, 58, 64, 30]: First, to quantify agreement in state space (attractor) geometry, we employed a Kullback-Leibler divergence defined across space, $D_{stsp}$ (see Appx. A.1 for details). For assessing agreement in long-term temporal properties, we used the Hellinger distance $D_{H}$ defined across power spectra of the true and model-generated trajectories. In addition, a short-term, $n$-step-ahead mean absolute prediction error (MAE) was evaluated. As has been repeatedly emphasized in the statistical and DSR literature [94, 48, 6, 58, 68, 30], prediction errors defined on time series are often only sensible on short time scales, because of the well known exponentially fast divergence of trajectories in chaotic DS (see Fig. 24).
Empirically observed time series are often just one-dimensional ($N^{*}=1$), but come from an inherently much higher-dimensional underlying DS. Our approach can in principle handle context signals of dimension $N^{*}<N$ smaller than those used in training by, e.g., just filling up $N-N^{*}$ dimensions in $\bm{C}$ with zeros. However, to work efficiently in a DSR setting, the context signal should represent the underlying DS as good as possible, for which it commonly needs to be lifted into a higher-dimensional space where sets explored by the dynamics ideally become diffeomorphic to those in the system’s true state space. Temporal delay embedding is the most popular technique to achieve this [80, 74]. For a $1d$ time series $\{x_{t}\}$, a $d$-dimensional embedding can be defined as [49]
@@ 4.1 Zero-shot DSR
Figs. 1 (top) & 3a present example zero-shot DSRs of new chaotic DS, which were not part of DynaMix’ training domain, from just a provided context signal $\bm{C}$ (see Figs. 11-15 for further examples). Despite the inherent challenge of predicting chaotic systems, where small differences in initial conditions lead to rapidly diverging trajectories (Fig. 24), DynaMix successfully captures the system’s temporal & geometrical behavior, i.e. accurately reconstructs the underlying state space and faithfully recovers the true system’s power spectrum. Fig. 3b highlights further aspects of DynaMix’ reconstructions: Even though not trained on any $2d$ systems, the model successfully recreates the limit cycle from the $2d$ Selkov system [77]. Notably, the model even generalizes to new initial conditions outside the scope of the context data, i.e. correctly infers properties of the state space beyond the short context trajectory (see Fig. 22 for further examples). Likewise, although trained only on $3d$ DS, DynaMix also successfully generalizes to higher-dimensional DS, as illustrated for the $6d$ Lorenz-96 system [54] in Fig. 34 (see Appx. A.6 for details on the setup and comparisons to other models). As further shown in Fig. 3c, the context length necessary to achieve good DSR quickly converges, with as few as $500$ time steps often sufficient. Fig. 3d illustrates that our approach also works across a surprisingly wide range (about an order of magnitude) of temporal resolutions at which the underlying DS was sampled, confirming its robustness w.r.t. sampling frequency.
We next compared the zero-shot, out-of-domain DSR performance of DynaMix to that of a number of recent TS foundation models, including various versions of Chronos [2], Mamba4Cast [5], TimesFM [20], and Tiny Time Mixers [26], using the long-term statistics $D_{stsp}$ and $D_{H}$ for evaluation (see Sect. A.1). As seen in Fig. 4a-b, DynaMix clearly and significantly outperforms all of these. In terms of short-term ($10$-step ahead) forecasts, for which TS foundation models (unlike DynaMix) are optimized, DynaMix performs on par with the (strongest in our batch) Chronos-t5 models, as shown in Fig. 4c. Table LABEL:tab:performance_DSR assembles further results on performance comparisons, while Figs. 16-21 provide specific examples. Moreover, a significant limitation of particularly the large Chronos models with millions of parameters is their extreme inefficiency in both inference time and computational costs, as compared in Fig. 5. In fact, their inference time can even exceed that of a custom-trained model. In contrast, DynaMix provides a highly efficient alternative, offering superior DSR performance while using orders of magnitude less parameters ($\approx 10$k in total) and computation time.
@@
To put DynaMix’ performance further in context, we also compared it to a couple of custom-trained DSR models, i.e. SOTA DSR models explicitly trained on the context data, incl. the plain AL-RNN trained by STF [7], Neural ODEs [37, 17], and reservoir computers which are commonly used for DSR [70] (see Appx. A.7 for details). Despite the custom trained models’ unduly advantage, performing – unlike DynaMix – only in-domain generalization [37], DynaMix keeps up with, and in some cases even outperforms, them (Appx. Tables 68).

### v3

@@ 1 Introduction
Figure 1: DynaMix achieves zero-shot DSR of attractor geometry and long-term temporal properties (power spectrum) from a short context signal while Chronos Ansari et al. (2024) fails.
Most real-world processes, from atmospheric phenomena and stock markets to brain activity or ecological networks, can be described as dynamical systems (DS) Mandelbrot and Hudson (2007); Buzsaki (2006); Tziperman et al. (1997); Mumby et al. (2007). Reconstructing these from observational data, called dynamical systems reconstruction (DSR), has been a long-standing challenge in scientific modeling Yu and Wang (2024); Durstewitz et al. (2023); Gilpin (2024). DSR goes beyond conventional time series (TS) modeling, as we wish to have a generative model of the underlying process which exhibits long-term behavior with the same temporal and geometrical signatures (Fig. 1), i.e. the same invariant or ‘climate’ statistics Patel and Ott (2023); Platt et al. (2023), as important in scientific applications. Achieving this usually requires special control-theoretic training techniques Mikhaeil et al. (2022); Hess et al. (2023) or loss objectives Platt et al. (2022, 2023); Jiang et al. (2023); Schiff et al. (2024) which accentuate the system’s long-term dynamics. Numerous deep learning approaches for DSR, based on recurrent neural networks (RNNs; Trischler and D’Eleuterio (2016); Durstewitz (2017); Vlachas et al. (2018); Cestnik and Abel (2019); Brenner et al. (2022); Rusch et al. (2022); Hess et al. (2023); Brenner et al. (2024a); Platt et al. (2023)), neural ODEs Chen et al. (2018); Karlsson and Svanström (2019); Alvarez et al. (2020); Ko et al. (2023), Koopman operators Brunton et al. (2021); Lusch et al. (2018); Otto and Rowley (2019); Azencot et al. (2020); Naiman and Azencot (2021); Geneva and Zabaras (2022); Wang et al. (2022), or library-based methods Brunton et al. (2016); Loiseau and Brunton (2018); Kaiser et al. (2018); Cortiella et al. (2021); Messenger and Bortz (2021), have been advanced over the years. However, all of these require purpose-training on the specific system observed and struggle to generalize beyond their training distribution Göring et al. (2024b).
Inspired by the strong in-context and zero-shot generalization abilities of LLMs Brown et al. (2020); Garg et al. (2022); Dong et al. (2022); Coda-Forno et al. (2023), there has been a push recently to develop models with likewise properties for the time series domain. Time series foundation models like Chronos Ansari et al. (2024, 2025) or Mamba4Cast Bhethanabhotla et al. (2024) are pre-trained on a huge database of real-world and artificial time series, and then tasked to forecast novel time series from which snippets are presented in-context, without any parameter fine-tuning. They are not built for DSR, however, and – as we show here – typically fail to properly capture a system’s long-term behavior and the structure of its attractors (Fig. 1; Zhang and Gilpin (2025)).
@@
  • Accurate zero-shot DSR: DynaMix achieves strong generalization across diverse DS, eliminating the need for fine-tuning while maintaining accuracy in reproducing attractor geometry and long-term statistics. No other model tested achieved this.
@@
  • Computational and parameter efficiency: The model reaches high performance with a very lightweight architecture ($\approx 10$k parameters) and small training corpus ($34$ systems), enabling orders of magnitude faster inference than other foundation models.
@@ 2 Related work
In DSR we seek to learn generative models from time series data that represent the underlying system dynamics with all its topological, geometrical, and temporal properties Durstewitz et al. (2023); Göring et al. (2024b); Platt et al. (2023); Gilpin (2023); Brunton et al. (2017); Brunton and Kutz (2019), in that sense providing an approximation to the underlying system’s governing equations. By definition, such a model should not only render viable short-term forecasts, but also reproduce the long-term evolution of the DS it has been trained on, both in state space and in the time domain. Methods approaching this goal have been founded on predefined function libraries, such as Sparse Identification of Nonlinear Dynamics (SINDy; Brunton et al. (2016); Loiseau and Brunton (2018); Cortiella et al. (2021); Kaiser et al. (2018); Messenger and Bortz (2021)), on reservoir computers Pathak et al. (2017); Platt et al. (2022, 2023), neural ODEs Chen et al. (2018); Karlsson and Svanström (2019); Alvarez et al. (2020); Ko et al. (2023), Koopman operators Brunton et al. (2021); Lusch et al. (2018); Otto and Rowley (2019); Azencot et al. (2020); Naiman and Azencot (2021); Geneva and Zabaras (2022); Wang et al. (2022), or different types of RNNs Trischler and D’Eleuterio (2016); Durstewitz (2017); Vlachas et al. (2018); Cestnik and Abel (2019); Brenner et al. (2022); Rusch et al. (2022); Hess et al. (2023); Brenner et al. (2024a). More important than the architecture itself seems to be their proper training in order to ensure long-term (invariant) statistics of the underlying system are met: Methods like sparse Mikhaeil et al. (2022); Brenner et al. (2022) or generalized Hess et al. (2023) teacher forcing allow RNN-generated trajectories to ‘explore the future’ whilst training, yet keep loss gradients in check. Alternatively, long-term statistics based on the observed system’s Lyapunov spectrum, fractal geometry, or invariant measures may be added to regularize the loss Platt et al. (2022, 2023); Jiang et al. (2023); Schiff et al. (2024), but require to compute these properties from the data first. Despite these advances, out-of-domain generalization remains a key challenge in DSR Göring et al. (2024b). Meta- and hierarchical learning models, trained across many DS simultaneously, have recently been designed as steps toward foundation models for DSR Brenner et al. (2025); Nzoyem et al. (2025); Kirchmeyer et al. (2022); Yin et al. (2021), but still require parameter fine-tuning and lack in-context inference capabilities.
@@
Large language models (LLMs) exhibit an impressive ability to infer patterns from prompts (context) and generalize to novel situations without retraining Brown et al. (2020); Garg et al. (2022); Dong et al. (2022); Coda-Forno et al. (2023), although the underlying reasons for this are still a matter of debate Lin and Lee (2024); Von Oswald et al. (2023); Xie et al. (2021). This inspired the development of general-purpose time series (TS) foundation models which could accurately forecast time series from a short segment provided ‘in-context’ Ansari et al. (2024); Ekambaram et al. (2024), without the need of task-specific fine-tuning Das et al. (2024). One idea is to simply use pretrained LLMs directly as TS forecasters Gruver et al. (2023); Rasul et al. (2023); Sun et al. (2023); Williams et al. (2024); Tang et al. (2025), but the inherent differences between textual and many temporal data, such as continuity, present significant challenges Tan et al. (2024). To overcome these, transformer-based architectures, similar in design to LLMs, such as Chronos Ansari et al. (2024, 2025) or TimesFM Das et al. (2024), were specifically pretrained on a large corpus of time series data. Promising zero-shot forecasting capabilities have also been achieved with alternative architectural designs, such as Tiny Time Mixers Ekambaram et al. (2024) or state-space models like Mamba Bhethanabhotla et al. (2024).
The success of TS foundation models raised hope these could also be utilized for zero-shot DSR, but the – to our knowledge – so far only previous study on this, based on Chronos, had mixed outcomes and fell short of a full DSR evaluation Zhang and Gilpin (2025): Successful DSR, if present at all, heavily depended on the initial conditions, and mechanisms such as context parroting visually gave the wrong illusion that features of the dynamics had been captured. In fact, here we show that existing TS foundation models are generally not capable of producing valid DSRs (Fig. 1), and highlight some of their failure modes. This is in contrast to the zero-shot DSR foundation architecture we develop here, based on models Brenner et al. (2024a) and training algorithms Mikhaeil et al. (2022) successful in DSR.
@@ 3.1 Model architecture
To enable zero-shot reconstruction of novel DS, we develop a specific mixture-of-experts (MoE) architecture that can be pretrained across many diverse DS (Fig. 10), with different experts possibly specializing in different dynamical regimes. As a SOTA DSR base model for the experts, we leverage a recent parameter-friendly RNN, which allows for highly efficient DSR training and is designed to yield topologically parsimonious and interpretable representations of DS it is being trained on, the Almost-Linear RNN (AL-RNN; Brenner et al. (2024a)):
@@
$\bm{w}^{att}_{t}=\sigma\left(\frac{\left|\bm{C}-\left(\bm{Dz}_{t}+\bm{\epsilon}\right)\bm{1}_{T_{C}}^{\top}\right|^{\top}\bm{1}_{N}}{\tau_{\text{att}}}\right)\in\mathbb{R}^{T_{C}}\;,$ (3)
@@ 3.2 Model training
For training, a collection of just $34$ different chaotic and cyclic DS is used Gilpin (2022), see Appx. A.2 for full specification and Fig. 10 for illustration. From each, multivariate time series $\bm{X}\in\mathbb{R}^{N\times T}$ are simulated ($\approx 6\times 10^{5}$ in total) and then standardized dimension-wise (likewise for the test set), of which the first $T_{C}<T$ column entries are defined as the context signal $\bm{C}=\bm{X}_{1:T_{C}}$. The experts are initialized at $t_{0}=T_{C}-\Delta t+1$ and forward-iterated until time $T$, such that they overlap for a period $\Delta t$ with the context. This is to ensure that the model learns to optimally utilize the context for generalization. A MSE loss is then computed across ground truth, $\bm{X}_{T_{C}-\Delta t+1:T}$, and respective model-generated time series $\hat{\bm{X}}$. The model is trained by sparse teacher forcing (STF; Brenner et al. (2022); Mikhaeil et al. (2022)), a control-theoretic technique specifically designed for DSR. STF replaces part of the forward-iterated latent states $\bm{z}_{t}$ by data-inferred states $\hat{\bm{z}}_{t}$ at optimally chosen intervals $\tau$, to avoid exploding gradients even for chaotic systems while enabling the DSR model “to explore the future” (see Mikhaeil et al. (2022) for detailed theoretical motivation). As shown in Fig. 26, STF with optimal $\tau$ is essential for achieving good zero-shot DSR results. Note that STF is only used for training and turned off at test time.
@@ 3.3 DSR evaluation & time series forecasting
We used two established measures to assess DSR quality Koppe et al. (2019); Brenner et al. (2022); Hess et al. (2023); Wood (2010); Zhang and Gilpin (2025); Platt et al. (2023); Mikhaeil et al. (2022); Pals et al. (2024); Gilpin (2023): First, to quantify agreement in state space (attractor) geometry, we employed a Kullback-Leibler divergence defined across space, $D_{stsp}$ (see Appx. A.1 for details). For assessing agreement in long-term temporal properties, we used the Hellinger distance $D_{H}$ defined across power spectra of the true and model-generated trajectories. Crucially, to properly assess the long-term dynamics, both these measures were evaluated in the limit of large $T=10,000$. In addition, a short-term, $n$-step-ahead mean absolute prediction error (MAE) was evaluated. As has been repeatedly emphasized in the statistical and DSR literature Wood (2010); Koppe et al. (2019); Brenner et al. (2022); Mikhaeil et al. (2022); Platt et al. (2021); Gilpin (2023), prediction errors defined on time series are often only sensible on short time scales, because of the well known exponentially fast divergence of trajectories in chaotic DS (see Fig. 24).
Empirically observed time series are often just one-dimensional ($N^{*}=1$), but come from an inherently much higher-dimensional underlying DS. To work efficiently in a DSR setting, the context signal should represent the underlying DS as good as possible, for which it commonly needs to be lifted into a higher-dimensional space where sets explored by the dynamics ideally become diffeomorphic to those in the system’s true state space. Temporal delay embedding is the most popular technique to achieve this Takens (1981); Sauer et al. (1991). For a $1d$ time series $\{x_{t}\}$, a $d$-dimensional embedding can be defined as Kraemer et al. (2021)
@@ 4.1 Zero-shot DSR
Figs. 1 (top) & 3a present example zero-shot DSRs of new chaotic DS, which were not part of DynaMix’ training domain, from just a provided context signal $\bm{C}$ (see Figs. 11-15 for further examples). Despite the inherent challenge of predicting chaotic systems, where small differences in initial conditions lead to rapidly diverging trajectories (Fig. 24), DynaMix successfully captures the system’s temporal & geometrical behavior, i.e. accurately reconstructs the underlying state space and faithfully recovers the true system’s power spectrum. Fig. 3b highlights further aspects of DynaMix’ reconstructions: Even though not trained on any $2d$ systems, the model successfully recreates the limit cycle from the $2d$ Selkov system SEL’KOV (1968). Notably, the model even generalizes to new initial conditions outside the scope of the context data, i.e. correctly infers properties of the state space beyond the short context trajectory (see Fig. 22 for further examples). Likewise, although trained only on $3d$ DS, DynaMix also successfully generalizes to higher-dimensional DS, as illustrated for the $6d$ Lorenz-96 system Lorenz (1996) in Fig. 34 (see Appx. A.6 for details on the setup and comparisons to other models). As further shown in Fig. 3c, the context length necessary to achieve good DSR quickly converges, with as few as $500$ time steps often sufficient. Fig. 3d illustrates that our approach also works across a surprisingly wide range (about an order of magnitude) of temporal resolutions at which the underlying DS was sampled, confirming its robustness w.r.t. sampling frequency.
We next compared the zero-shot, out-of-domain DSR performance of DynaMix to that of a number of recent TS foundation models, including various versions of Chronos Ansari et al. (2024), Chronos-2 Ansari et al. (2025), Panda Lai et al. (2025), Mamba4Cast Bhethanabhotla et al. (2024), TimesFM Das et al. (2024), and Tiny Time Mixers Ekambaram et al. (2024), using the long-term statistics $D_{stsp}$ and $D_{H}$ for evaluation (see Sect. A.1). As seen in Fig. 4a-b, DynaMix clearly and significantly outperforms all of these. In terms of short-term ($10$-step ahead) forecasts, for which TS foundation models (unlike DynaMix) are optimized, DynaMix performs on par with the strongest competitors in our batch, as shown in Fig. 4c. This is especially surprising when compared to Panda, which included variations of our test set DS in its training repertoire, biasing the evaluation in its favor. Table LABEL:tab:performance_DSR assembles further results on performance comparisons, while Figs. 16-21 provide specific examples. Moreover, a significant limitation of particularly the large Chronos models with millions of parameters is their extreme inefficiency in both inference time and computational costs, as compared in Fig. 5. In fact, their inference time can even exceed that of a custom-trained model. In contrast, DynaMix provides a highly efficient alternative, offering superior DSR performance while using orders of magnitude less parameters ($\approx 10$k in total) and computation time.
@@
To put DynaMix’ performance further in context, we also compared it to a couple of custom-trained DSR models, i.e. SOTA DSR models explicitly trained on the context data, incl. the plain AL-RNN trained by STF Brenner et al. (2024a), Neural ODEs Göring et al. (2024a); Chen et al. (2018), and reservoir computers which are commonly used for DSR Platt et al. (2023) (see Appx. A.7 for details). Despite the custom trained models’ unduly advantage, performing – unlike DynaMix – only in-domain generalization Göring et al. (2024a), DynaMix keeps up with, and in some cases even outperforms, them (Appx. Tables 68).